# Improving the solver for the Balitsky-Kovchegov evolution equation with Automatic Differentiation

### v1

@@ Appendix A Implementation details
In this Section we provide some additional details concerning the implementation.
Listing 4: Wrapper of the $\exp$ function.
@@
Listing 5: Excerpt from the dipole class definition.
template <typename T> class dipole{
public:
dual<T> *amplitude;
int dipole_size;
explicit dipole(int range){
dipole_size=range;
amplitude = (dual<T>*) malloc(sizeof(dual<T>)*range);
inline void set(dual<T> x, int n){
amplitude[n].val = x.val;
amplitude[n].derr = x.derr;
amplitude[n].der2r = x.der2r;
for(int i=0; i<NN; i++)
amplitude[n].der[i] = x.der[i];
for(int i=0; i<NN_hessian; i++)
amplitude[n].der2[i] = x.der2[i];
inline void set(T x, int n){
amplitude[n].val = x;
amplitude[n].derr = 0.0;
amplitude[n].der2r = 0.0;
for(int i=0; i<NN; i++)
amplitude[n].der[i] = 0.0;
for(int i=0; i<NN_hessian; i++)
amplitude[n].der2[i] = 0.0;
inline dual<T> get(int n){
if(n < dipole_size)
return amplitude[n];
else
return amplitude[dipole_size-1];
//function that returns the dipole amplitude at position x
//s is the log r increment on the logarithmic scale
//intmax is the size of the array
inline pair<T> get(dual<T> x, T s, int intmax){
dual<T> z;
//go from linear scale to logarithmic scale
T loggx = log(x.val);
//find the nearest nodes in the grid
int site_c = ceil((loggx-Slogmin)/s);
int site_f = floor((loggx-Slogmin)/s);
//deal with boundary cases
if( site_c > intmax-1 )
site_c = intmax-1;
if( site_f < 0 )
site_f = 0;
if( site_f > intmax-1 )
site_f = intmax-1;
if( site_c < 0 )
site_c = 0;
if(site_c != site_f){
//linear interpolation
dual<T> val_c = amplitude[site_c];
dual<T> val_f = amplitude[site_f];
dual<T> b = (val_c - val_f)
* ( (T) (1.0 / (1.0*(site_c - site_f))));
dual<T> a = val_c - (val_c - val_f)
* ( (T)((1.0 * site_c) / (1.0*(site_c - site_f))));
dual<T> z_tmp = a + b * ((T)(loggx-Slogmin)/s);
z.val = z_tmp.val;
//(S[f[x]])’ = S’[f[x]] * f’[x]
z.derr = z_tmp.derr * x.derr;
//(S[f])’ = S’[f]
for(int i=0; i<NN; i++){
z.der[i] = z_tmp.der[i];
//(S[f[x]])” = S’[f[x]] * f”[x] + S”[f[x]] * (f’[x])^2
z.der2r = z_tmp.derr * x.der2r
+ z_tmp.der2r * x.derr * x.derr;
//(S[f])” = S”[f]
for(int i=0; i<NN_hessian; i++){
z.der2[i] = z_tmp.der2[i];
}else{
z.val = amplitude[site_c].val;
z.derr = amplitude[site_c].derr * x.derr;
z.der2r = amplitude[site_c].derr * x.der2r
+ x.derr * x.derr * amplitude[site_c].der2r;
for(int i=0; i<NN; i++)
z.der[i] = amplitude[site_c].der[i];
for(int i=0; i<NN_hessian; i++){
z.der2[i] = amplitude[site_c].der2[i];
return z;
@@ Appendix B TMD in Color Glass Condensate
In high-energy limit strong intersections can be described within effective field theory called Color Glass Condensate (CGC) [50, 51, 52, 53, 54, 55, 56]. In this picture, the parton (gluon or quark) with high longitudinal momentum scatters off the classical gluonic field $A^{a}_{\rm cl}(z^{-},\bm{x}_{\perp})$. Degrees of freedom are given by Wilson lines:
@@
where only $+$ component of the field contributes.

### v2

@@ Appendix A Implementation details
In this Section, we provide some additional, and technical details concerning the implementation.
In the listing 1 we show the class which implements the dual numbers. It contains one parameter adjustable by the user at compile time, namely $NN$ that defines the number of differentiable parameters to be included in the calculation. This number does not include the radial variable which is kept separate. The class contains containers where the first and second derivatives are stored and accumulated.
//number of parameters with derivatives
const int NN = 2;
//number of independent entries in the hessian matrix
const int NN_hessian = NN*NN - NN*(NN-1)/2;
template <typename T> class dual {
public:
T val;
T derr;
T der2r;
T der[NN];
T der2[NN_hessian];
dual(T v){
val = v;
derr = 0.0;
der2r = 0.0;
for(int i=0; i<NN; i++)
der[i] = 0.0;
for(int i=0; i<NN_hessian; i++)
der2[i] = 0.0;
Listing 2 shows the overloading of the multiplication operator for dual numbers. Infinitesimal arithmetics for the first and second derivatives are implemented separately for the radial variable and for the set of differentiable parameters. Similar overloading is done for all other basic operations appearing in the algorithm. Additional functions are added to treat mixed-type operations, for instance, the multiplication of a dual number by a real number.
template <typename T> dual<T> operator*(dual<T> a, dual<T> b){
dual<T> c;
c.val = a.val*b.val;
c.derr = a.derr*b.val + a.val*b.derr;
c.der2r = a.der2r*b.val + 2.0*a.derr*b.derr + a.val*b.der2r;
for(int i=0; i<NN; i++)
c.der[i] = a.der[i]*b.val + a.val*b.der[i];
for(int i=0; i<NN; i++)
for(int j=i; j<NN; j++){
int index = i*NN+j-i*(i+1)/2;
c.der2[index] = a.der2[index]*b.val
+ a.der[i]*b.der[j] + a.der[j]*b.der[i]
+ a.val*b.der2[index];
return c;
Listing 3 contains a wrapper of the exponential function. Apart from the single call to the standard library $\exp$ function, the calculation of the derivatives is implemented.
Listing 3: Wrapper of the $\exp$ function.
@@ Appendix B TMD in Color Glass Condensate
In high-energy limit, strong intersections can be described within effective field theory called Color Glass Condensate (CGC) [52, 53, 54, 55, 56, 57, 58]. In this picture, the parton (gluon or quark) with high longitudinal momentum scatters off the classical gluonic field $A^{a}_{\rm cl}(z^{-},\bm{x}_{\perp})$. Degrees of freedom are given by Wilson lines:
@@
where only ++ component of the field contributes.